1. A ball is thrown off a cliff overlooking a lake. The cliff is 9.60 m above the lake. The ball has an initial speed of 2.85 m/s and is thrown at an angle of 20.0° below the horizontal. (a) Calculate the time after the ball is thrown until it strikes the water. (b) Calculate the ball’s speed just before striking the water. (c) How far from the base of the cliff does the ball hit the lake? (d) Sketch the velocity versus time graphs for the ball from the time it was launched until it hits the ground. You need to draw two graphs, one for the horizontal velocity and one for the vertical velocity. Numerical labels are not required.
Answer #1

a)

motion along vertical

h = v sin x * t - 0.5 gt^2

9.6 = 2.85 sin 20 * t - 4.9 t^2

solving for t

t = 1.306 s

======

b)

using 3rd equation of motion

v'^2 = v^2 + 2 g h

v' ^2 = 2.85^2 + 2* 9.8* 9.6

v' = 14.01 m/s

======

c)

horizontal distanve

D = 2.85 * cos 20 * t

D = 3.4976 m

======

Comment before rate in case any doubt, will reply for sure.. goodluck